OpenStudy (anonymous):

two truck drivers leave a cafe at the same time traveling in opposite directions. one truck goes 7 mph faster than the other one. after 4 hours they are 404 miles apart. how fast is each truck going?

OpenStudy (lgbasallote):

let x = speed of truck A x + 7 = truck B <--because 7 mph faster after 4 hours the distance of truck B would be 4(x+7) according to distance = speed x time after 4 hours the distance of truck A would be 4x according to distance = speed x time it says that the distance between them is 404 so let's add 4x + 4(x+7) = 404 4x + 4x + 28 = 404 8x = 404 - 28 8x = 376 solve for x

OpenStudy (lgbasallote):

do you get that?

OpenStudy (anonymous):

yes i get it. thanks

OpenStudy (lgbasallote):

<tips hat>